Using Elimination Method solve
x+8y=41
-x+9y=27
find the ordered pair

Found 2 solutions by Alan3354, solver91311:
Answer by Alan3354(69443)

x +8y=41
-x+9y=27
================ Add

Answer by solver91311(24713)



Add the two equations term by term to eliminate the variable. Solve for . Substitute the value of into either original equation and then solve for . Create the ordered pair

John

My calculator said it, I believe it, that settles it